OpenStudy (howard-wolowitz):

Can someone explain how to get the answer in this problem: Yes, these are separate questions!

OpenStudy (thomas5267):

Okay. First note that \(\left((f^{-1})^{-1}\right)(x)=f(x)\). (i.e. the inverse function of the inverse function is the function itself (there are some restrictions to it but you don't have to worry about it now)) Secondly, option 3 and 4 are not correct because the inverse of a straight line is a straight line. For example, option 3 is the inverse of \(f(x)=2x-3\). So it is either option 1 or 2. Try to invert both option 1 and 2 to yield the original function and see if it fits the graph.

OpenStudy (howard-wolowitz):

I'm graphing it now

OpenStudy (howard-wolowitz):



OpenStudy (howard-wolowitz):

us that right for A

OpenStudy (thomas5267):

\[ x=\log_2(y+3) \] Rearrange it to y=sth.

OpenStudy (howard-wolowitz):

what do you mean

OpenStudy (thomas5267):

Do you know how to invert a function?

OpenStudy (howard-wolowitz):

i guess not

OpenStudy (thomas5267):

Let say you want to invert \(f^{-1}(x)=\dfrac{x-3}{2}\). First write \(y=\dfrac{x-3}{2}\). Then you swap the x and y, i.e. \(x=\dfrac{y-3}{2}\). You then rearrange the equation such that y is only on one side and x is only on the other side. That is:\[ \begin{align*} x&=\frac{y-3}{2}\\ 2x&=y-3\\ y&=2x+3 \end{align*} \] Try invert the third option.

OpenStudy (howard-wolowitz):

y=2x-3

OpenStudy (thomas5267):

How about the first option?

OpenStudy (howard-wolowitz):

umm

OpenStudy (howard-wolowitz):

I dont know how to do that one

OpenStudy (thomas5267):

I will do the second one for you. \[ \begin{align*} y&=\log_2(x)+3\\ x&=\log_2(y)+3\\ x-3&=\log_2(y)\\ y&=2^{x-3}\text{ as }2^{\log_2(y)}=y.\quad \text{In fact, }n^{\log_n(x)}=x \end{align*} \] Try do something similar.

OpenStudy (howard-wolowitz):